[Brahms-l] Shift report 20050130 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jan 30 2005 - 07:42:28 EST
1. Very nice store since 2am plus the end of the previous one gives almost 6 hours of beam during this shift.
2. In the short period between the two stores I had Magnet Crew changed the tabs on D2 so that we can go to full field. Hence the following runs has been recorded:
12674 - FS @ 8 deg @ A1723  - 49k tr2.
12675 - D1 at A861, rest at 0 - 76k tr6.
12676 - D2 at A723, rest at 0 - 93k tr6.
12677 and 12678 - pedestal runs.
12679 - FS @ 8 deg @ A3450 -  81k tr2.
12680 - FS @ 8 deg @ A3450 - 107k tr2.
12681 - FS @ 8 deg @ A3450 -  25k tr2.
12682 - FS @ 8 deg @ A3450 -  61k tr2.
12683 and 12684 - pedestal runs.
3. MRS1440 still does not work. None other detector problems seen.


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jan 30 07:42:46 2005

This archive was generated by hypermail 2.1.8 : Sun Jan 30 2005 - 07:43:08 EST